# Linear Discriminant Regularized Regression

### v1

@@ B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$
$\quad+\left({\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
@@
$\widehat{\Sigma}$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
@@
The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 12. We bound the second term by invoking 1 and Lemma 11 with $u={\mathbb{M}}^{\top}v_{1}$ and $v={\mathbb{M}}^{\top}v_{2}$. This yields
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}\over n}$
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$ (B.11)
@@
and after we invoke Lemma 10 twice with $v=v_{1},u={\mathbb{M}}^{\top}v_{2}$ and $v=v_{2},u={\mathbb{M}}^{\top}v_{1}$, respectively, we get that
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (B.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (B.13)
@@
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 13 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (B.5), (B.6) and (B.7) of Lemma 9 gives that, with probability at least $1-n^{-2}$,
$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\quad+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
@@
The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim. Regarding the operator norm bound, a standard discretization argument gives
@@
$\|B^{\top}{\mathbb{M}}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
$\|{\mathbb{M}}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
@@
$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{\mathbb{M}}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
$\quad+\|{\mathbb{M}}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
@@
Appendix C Auxiliary lemmas

### v2

@@ C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$
$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
@@
$\widehat{\Sigma}$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
@@
The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 17. We bound the second term by invoking 1 and Lemma 16 with $u={M}^{\top}v_{1}$ and $v={M}^{\top}v_{2}$. This yields
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}\over n}$
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (C.11)
@@
and after we invoke Lemma 15 twice with $v=v_{1},u={M}^{\top}v_{2}$ and $v=v_{2},u={M}^{\top}v_{1}$, respectively, we get that
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (C.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (C.13)
@@
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (C.5), (C.6) and (C.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,
$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{M}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\quad+\|{M}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
@@
The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim.
Regarding the operator norm bound, a standard discretization argument gives
@@
$\|B^{\top}{M}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
$\|{M}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
@@
$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{M}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{M}{M}^{\top}Bu}\sqrt{\log(n)\over n}$
$\quad+\|{M}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{M}{M}^{\top}Bu}\sqrt{\log(n)\over n}$
@@
The following lemma provides upper bounds of the operator norm of $(\widehat{\Sigma}-\Sigma)$.
Lemma 20.
Under model (1.1) and 3, assume $(p+L)\log(n)\leq n$. Then with probability at least $1-n^{-2}$, the following holds uniformly over $v\in\mathcal{S}^{p}$
$v^{\top}(\widehat{\Sigma}-\Sigma)v~{}\lesssim~{}v^{\top}\Sigma v\left(\sqrt{p\log(n)\over n}+\sqrt{L\log(n)\over n}\right).$
As a result, with the same probability, we have
$\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma).$
Proof.
The proof of the first statement follows from the same arguments of proving Lemma 12 and is thus omitted.
For the second statement, for any $v\in\mathcal{S}^{p}$, we have
$v^{\top}\widehat{\Sigma}v=v^{\top}\Sigma^{1/2}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right)\Sigma^{1/2}v\geq v^{\top}\Sigma v~{}\lambda_{p}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right).$
Since, by Weyl’s inequality and on the event that the first result holds,
$\lambda_{p}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right)\geq 1-\|\Sigma^{-1/2}(\widehat{\Sigma}-\Sigma)\Sigma^{-1/2}\|_{\rm op}\geq 1-\left(\sqrt{p\log(n)\over n}+\sqrt{L\log(n)\over n}\right)$
we conclude
$v^{\top}\widehat{\Sigma}v\geq{1\over 2}v^{\top}\Sigma v$
uniformly over $\mathcal{S}^{p}$, completing the proof. ∎
Appendix D Auxiliary lemmas